Michael wants to receive a total of $760 on his investment. He wants to invest part of his $8,000 in a bond that pays 12% annual interest and the rest in a savings account that pays 8%. How much should he invest at each rate?
Thanks for any help you can provide - I'm stuck!

Answer by Earlsdon(5032)

Let x = the amount invested at 12%, then (8000-x) = the amount invested at 8%
and the interest earned on this is to be $760.00 You can write the equation for this as follows:
(12%)x + (8%)($8000-x) = $760 Simplfy and solve for x. Change the percentages to decimals.
0.12x + 0.08($8000-x) = $760
0.12x + $640 - 0.08x = $760
0.04x + $640 = $760
0.04x = $120
x = $3,000 This is the amount Michael should invest at 12%
$8,000-x = $5,000 This is the amount Michael should invest at 8%